OpenStudy (anonymous):

Systems & Equations: Best Rentals charges a daily fee plus a mileage fee for renting its cars. Barney was charged $ 123.00 for 3 days and 300 miles, while Mary was charged $ 216.00 for 5 days and 600 miles. What does Best Rental charge per day, for mileage?

OpenStudy (gw2011):

Let x=daily fee Let y=mileage fee 3x+300y=123 Multiply by 5 to eliminate x 5x+600y=216 Multiply by -3 to eliminate x 15x+1500y=615 -15x-1800y=-648 -300y=-33 y=-33/-300=0.11 y=11 cents per mile 3x+300(0.11)=123 3x+33=123 3x=90 x=$30 daily fee Therefore, Best Rental charges a daily fee of $30 and 11 cents mileage fee

OpenStudy (anonymous):

Thanks.